When a certain wire is connected across a voltage of 1 V, the resulting current through the wire is 1 A. If we connect the same wire across a voltage of 2 V, the current will be

A 2A

B 4A

C 0.25A

D 1A

E 0.5A
Answer #1
Concepts and reason

The concept used to solve this problem is Ohm’s law.

Use Ohm’s law to find the current through the wire.

Fundamentals

Ohm’s law states that, “the potential difference across the conductor is proportional to the current through the conductor”.

The expression for the voltage is as follows:

V=IRV = IR

Here, the current is II, the voltage is VV, and the resistance is RR.

The incorrect options for current through the wire when the wire is connected across the voltage of 2V2\,{\rm{V}}are given below:

• B) 4A4\,{\rm{A}}

• C) 0.25A0.25\,{\rm{A}}

• D) 1A1\,{\rm{A}}

• E) 0.5A0.5\,{\rm{A}}

The expression for the voltage is as follows:

V=IRV = IR …… (1)

Rearrange the above expression for RR.

R=VIR = \frac{V}{I} …… (2)

When the wire is connected across the voltage of 1V1\,{\rm{V}}, then the resistance of the wire is calculated as follows:

Substitute 1V1\,{\rm{V}} for VV and 1A1\,{\rm{A}} for II in equation (2).

R=1V1A=1Ω\begin{array}{c}\\R = \frac{{1\,{\rm{V}}}}{{1\,{\rm{A}}}}\\\\ = 1\,\Omega \\\end{array}

Rearrange the expression (1) for II.

I=VRI = \frac{V}{R}

When the wire is connected across the voltage of 2V2\,{\rm{V}}, then the current through the wire is calculated as follows:

Substitute 2V2\,{\rm{V}} for VV and 1Ω1\,\Omega for RR in equation (2).

I=2V1Ω=2A\begin{array}{c}\\I = \frac{{2\,{\rm{V}}}}{{1\,\Omega }}\\\\ = 2\,{\rm{A}}\\\end{array}

The correct option for the current through the wire when the wire is connected across the voltage of 2V2\,{\rm{V}} is given below:

• A) 2A2\,{\rm{A}}

Therefore, the current through the wire when it is connected across the voltage of 2V2\,{\rm{V}}is 2A2\,{\rm{A}}.

Ans:

Thus, the current through the wire when it is connected across the voltage of 2V{\bf{2}}\,{\bf{V}}is 2A{\bf{2}}\,{\bf{A}}.